Don left Vancouver and drove towards Hope BC at a speed of 120km/h. Brad left Hope at the same time and drove towards Vancouver at a speed of 90km/h. If hope is 122km from Vancouver how far had they driven when they passed eachother on the highway?
Found 2 solutions by mananth, greenestamps:
Answer by mananth(16946)

Let x be the distance they meet from Vancouver
Don travels x km in time t
Brad travels 122-x km in time t
t = d/s
t(don) = x/120
t(brad) = (122-x)/90
t(brad) = t(Don)
x/120 = (122-x)/90
simplify
x/4 = (122-x)/3
3x= 4(122-x)
3x= 488-4x
7x = 488
x=69.7

Answer by greenestamps(13206)



You didn't ask how far EACH of them had traveled when they passed each other....

So the answer is trivial: if Hope is 122 km from Vancouver, then they (TOGETHER!) had driven 122 km when they passed each other.

If you wanted to know how far EACH had traveled, then note that the ratio of their speeds is 120:90 = 4:3; that means that when they passed each other Don had driven 4/7 of the 122 km and Brad had driven 3/7 of the 122 km.